A model turbine having diameter 0.4 m develops 15 kW when working under a head of 5m at

its best speed of 1600 rpm. If a geometrical similar turbine having diameter 1.6m operates at

the same efficiency under 18 m head. Find the speed and power developed by the actual turbine.


1
Expert's answer
2021-03-16T08:41:37-0400

"\dfrac{H_1}{(N_1D_1)²}= \dfrac{H_2}{(N_2D_2)²}"


"\dfrac{5}{(1600× 0.4)²}= \dfrac{18}{(1.6 N_2)²}"


"N_2 = 758.95" rpm


Equating the power coefficients;

"\dfrac{P_1}{ρ_1N_1^3 D_1^5}= \dfrac{P_2}{ρ_2N_2^3 D_2^5}"


Considering the fluids to be incompressible, and same for both the prototype and model,

we have,


"\dfrac{P_1}{N_1^3 D_1^5}= \dfrac{P_2}{N_2^3 D_2^5}"


"\dfrac{15}{1600^3 × 0.4^5}= \dfrac{P_2}{758.95^3× 1.6^5}"


"3.58 \u00d7 10^{-7} = 2.18 \u00d7 10^{-10} \u00d7 P_2"


"P_2 =" 1639.35 kW


"P_2 = 1.64\ MW"